OpenStudy (tylermckinney16):

Jose worked 36 hours last week. Since his hourly pay rate is $15.25 per hour, what is his gross pay?

OpenStudy (tylermckinney16):

@tigerlover

OpenStudy (3mar):

his gross pay=36*15.25=549 $

OpenStudy (3mar):

Hope that helps

tigerlover (tigerlover):

Tbh idk anything about job payments yet. sorry! :)

OpenStudy (rolldextro):

Your gross pay is the amount of money you would make before any adjustments. So in order to find the amount of money you made you would do the following \[G.P=36 hr* {\frac{15.25$}{1 hr} }=549$\]